OpenStudy (anonymous):

5a = 6-b 5a = 4-b Using the substitution method this linear equation ..I am saying that it has one solution? Would this be correct?

hero (hero):

No solution

hero (hero):

5a + b = 6 5a + b = 4 Impossible

OpenStudy (wildryc):

Not really? 6-b = 5a = 4-b, correct? a =(6-b)/5 and a= (4-b)/5

hero (hero):

Put it this way: If you subtracted the equations, you end up with 0 = 2 which is false.

OpenStudy (anonymous):

no solution